OpenStudy (anonymous):

An archer releases an arrow from a shoulder height of 1.39m. When the arrow hits the target 18m away, it hits point A (at 1.28m). When the target is removed, the arrow lands 45m away. Find the maximum height of the arrow along its parabolic path. Can someone PLEASE give a detailed description of how to solve this and show work? I've been sitting here for over an hour trying to figure this out.

OpenStudy (phi):

parabolic path means a parabola, which means an equation of the form y = ax^2 + b x + c here y will be the height, x is the distance from the archer. the max height will be the vertex of the parabola, which occurs at x= -b/(2a) so you have to find the coefficients a, b and c.

OpenStudy (phi):

An archer releases an arrow from a shoulder height of 1.39m we put the archer at x=0, so this point is (0,1.39) When the target is removed, the arrow lands 45m away. the height of the arrow on the ground is 0, so this point is (45,0) we need one more point.

OpenStudy (phi):

he arrow hits the target 18m away, it hits point A (at 1.28m). x= 18 y will be the height above the ground. But I don't know the height of point A. Is there a picture ?

OpenStudy (anonymous):

Wouldn't 1.28m be the height?

OpenStudy (phi):

maybe... it is not clear from the write-up. Is there a picture ?

OpenStudy (anonymous):

Yeah hold on!

OpenStudy (amistre64):

had this question the other day :)

OpenStudy (amistre64):

you are given 3 points to define a parabola with

OpenStudy (amistre64):

(0,shoulder) (18,A) (45,0)

OpenStudy (phi):

yes, the height is 1.28 so the third point is (18.1.28)

OpenStudy (phi):

use each point in the "generic equation" y = a x^2 + b x + c to generate 3 equations with 3 unknowns. can you do that ?

OpenStudy (amistre64):

well, we would already now one of the "unknowns" :) reducing it to 2 eqs in 2 unknowns

OpenStudy (anonymous):

I'm still really confused! How do I write the equations with the unkowns and what do I do with them?

OpenStudy (phi):

to start at the beginning: we are told we have a parabola (parabolic path) we (might? should?) know a parabola is y = a x^2 + b x + c where a , b and c are unknown coefficients to solve for 3 unknowns, we need 3 different equations. notice we have 3 (x,y) points. If we take an (x,y) point example: (0,1.39) and replace x and y with 0 and 1.39, we will get an equation with unknowns a,b and c can you do that ?

OpenStudy (anonymous):

So it would be like a system of equations thing?

OpenStudy (phi):

yes

OpenStudy (anonymous):

Okay hold on let me see if I can do this..

OpenStudy (anonymous):

Okay I'm really bad at 3 part systems of equations... I got these equations: 1.39=a(0)^2 + b(0) +c 0=a(45)^2 + b(45) + c 1.28 = a(18)2 + b(18) + c So then how do I find the values of the variables? Having 3 equations always messes me up!

OpenStudy (phi):

you can "solve" the first one easily.

OpenStudy (anonymous):

Oh yeah so wouldn't it be c=1.39?

OpenStudy (phi):

yes. so replace c with 1.39 in the other 2 equations. now it's 2 eq's and 2 unknowns

OpenStudy (anonymous):

So then would I do like the elimination method? Sorry I'm usually able to figure out these problems but today my mind has just gone blank...

OpenStudy (phi):

these are ugly numbers. but you now have 0=a(45)^2 + b(45) + 1.39 1.28 = a(18)2 + b(18) + 1.39 if we add -1.28 to both sides of the 2nd equation we get a(18)^2 + b(18) + 0.11=0 you now have \[ a(18)^2 + b(18) + 0.11=0 \\a(45)^2 + b(45) + 1.39=0 \]

OpenStudy (phi):

I would divide the first equation by 18, and the 2nd by 45 to get \[ 18a + b +\frac{0.11}{18} = 0 \\ 45a + b + \frac{1.39}{45} = 0\]

OpenStudy (anonymous):

Okay cool...what's the next step?

OpenStudy (anonymous):

Do we eliminate b?

OpenStudy (phi):

subtract the two equations

OpenStudy (anonymous):

Okay I got -27a - 0.0247 = 0?

OpenStudy (phi):

keep lots of decimals for this problem.... but now solve for a

OpenStudy (anonymous):

a = -9.15?

OpenStudy (phi):

a= 0.02477777/-27 that can be -9.15

OpenStudy (phi):

*can't

OpenStudy (anonymous):

Okay I got -9.17695185E-4 What does the E mean?

OpenStudy (phi):

scientific notation. E stands for "exponent" that number is \[ -9.17695185\cdot 10^{-4} = -0.000917695185\] but it is easier to keep it in scientific notation

OpenStudy (phi):

now solve for b using \[ 18a + b +\frac{0.11}{18} = 0 \]

OpenStudy (anonymous):

My calculator says "error" when I try to multiply 18 and -9.17695185 x 10^-4, so should I just use the decimal number instead??

OpenStudy (phi):

does your calculator have an "EXP" button ? if so, try 2 EXP 2 = you should see 200

OpenStudy (anonymous):

Yeah I found it!

OpenStudy (phi):

so type in -9.17695185 EXP -4 (and save it in memory if you know how)

OpenStudy (anonymous):

Okay I got b = -0.0104074023

OpenStudy (phi):

that looks good, except it should be positive

OpenStudy (anonymous):

Oh right sorry I didn't mean to put that negative there! Okay so now what do we do??

OpenStudy (phi):

Find the maximum height of the arrow along its parabolic path. that means find the vertex of the parabola see http://hotmath.com/hotmath_help/topics/vertex-of-a-parabola.html

OpenStudy (anonymous):

Okay so the vertex is (5.67, 1.42)

OpenStudy (anonymous):

So the max. height is 1.42m. Thank you SO much for your help @phi I couldn't have done it without you! <3

OpenStudy (phi):

yes, that looks good